# Inversion of adjunction for quotient singularities III: semi-invariant case

### 10 Dec 2023 v1

@@ Abstract.
We prove the precise inversion of adjunction formula for finite linear group quotients of complete intersection varieties defined by semi-invariant equations. As an application, we prove the semi-continuity of minimal log discrepancies for them. These results extend the results in our first paper, where we prove the same results for complete intersection varieties defined by “invariant equations”.
@@ 1. Introduction
The main purpose of this paper is to relax the assumptions of “invariant equations” in the results in [NS22] and extend them to the case of “semi-invariant equations”. For a finite subgroup $G\subset{\rm GL}_{N}(k)$, we say that $f\in k[x_{1},\ldots,x_{N}]$ is $G$-semi-invariant if for any $\alpha\in G$, there exists $s\in k^{\times}$ such that $\alpha(f)=sf$. In this paper, we show that the LSC conjecture is true for a finite linear group quotient of a complete intersection variety defined by semi-invariant equations. Note that this class of singularities includes all $3$-dimensional terminal singularities.
Theorem 1.2 ($=$ Theorem 7.4).
@@
Let $N$ be a positive integer. Set $R:=k[x_{1},\ldots,x_{N}]$. Let $d$ be a positive integer, and let $\xi\in k$ be a primitive $d$-th root of unity. Suppose that a finite group $G\subset\operatorname{GL}_{N}(k)$ of order $d$ acts on $\overline{A}=\operatorname{Spec}R$. Let $\overline{B}\subset\overline{A}$ be a complete intersection variety defined by $G$-semi-invariant elements $f_{1},\ldots,f_{c}\in(x_{1},\ldots,x_{N})\subset R$. We denote by $B:=\overline{B}/G$ its quotient variety. Since $G$ is a finite group, each $\gamma\in G$ can be diagonalized with some new basis $x_{1}^{(\gamma)},\ldots,x_{N}^{(\gamma)}$. Let $\operatorname{diag}\left(\xi^{e_{1}},\ldots,\xi^{e_{N}}\right)$ be the diagonal matrix with $10\leq e_{i}\leq d-$ 1. Then we define a $k$-algebra homomorphism ${\lambda}^{*}_{\gamma}:R\to R[t^{1/d}]$ as follows:
@@
In [NS22], we study the minimal log discrepancies of $B$ via the arc space of $\overline{B}^{(\gamma)}$ using Denef and Loeser’s theory in [DL02].
However, if $f_{i}$ is just assumed to be $G$-semi-invariant, $\lambda^{*}_{\gamma}(f_{i})$ is not necessarily an element of $R[t]$. Thus, the first difficulty is how to redefine $\overline{B}^{(\gamma)}$. For each $i$ and $\gamma\in G$, we define $w_{\gamma}(f_{i}):=\frac{a}{d}$ when $\gamma(f_{i})=\xi^{a}f_{i}$ with $10\leq a\leq d-$ 1. Then, we have $\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\in R[t]$. By using this invariant $w_{\gamma}$, we change the definition of $\overline{B}^{(\gamma)}$ as follows:
@@
Theorem 1.6 ($=$ Theorem 5.3).
Let $n:=N-c$. Let $x\in B$ be the image of the origin of $\overline{A}=\mathbb{A}_{k}^{N}$. Let $\mathfrak{a}\subset\mathcal{O}_{B}$ be a non-zero ideal sheaf and $\delta$ a positive real number. Then
$\operatorname{mld}_{x}(B,\mathfrak{a}^{\delta})$ $=\inf_{\gamma\in G,\ b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C_{\gamma,b_{1},b_{2}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\delta b_{1}\right\}$
$=\inf_{\gamma\in G,\ b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C^{\prime}_{\gamma,b_{1},b_{2}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\delta b_{1}\right\}$
@@
By the theorem of Ein, Mustaţă, and Yasuda ([EMY03]), the minimal log discrepancy $\operatorname{mld}_{x}(B,\mathfrak{a}^{\delta})$ is described by the codimensions of the corresponding contact loci in the arc space $B_{\infty}$ of $B$. Therefore, in order to prove Theorem 1.6, we need to compare the codimension of a cylinder $C\subset B_{\infty}$ and the codimension of the corresponding cylinder $C^{\prime}\subset\overline{B}^{(\gamma)}_{\infty}$. By the same argument in [NS22] (using the theory of Denef and Loeser in [DL02]), the difference can be measured by comparing $\Omega_{\overline{B}^{(\gamma)}/k[t]}^{n}$ and $\omega_{B}$ via the morphism $\overline{B}^{(\gamma)}\to B$.
@@
$L_{\overline{B}^{(\gamma)}}:=\bigl{(}\operatorname{det}^{-1}(I_{B}/I_{B}^{2})\bigr{)}\big{|}_{\overline{B}^{(\gamma)}}\otimes_{\mathcal{O}_{\overline{B}^{(\gamma)}}}\left(\Omega_{\overline{A}^{(\gamma)}/k[t]}^{N}\right)\big{|}_{\overline{B}^{(\gamma)}},$
@@
$B_{4}:=\overline{B}^{(\gamma)}\times_{\operatorname{Spec}k[t]}\operatorname{Spec}k[t^{1/d}]\longrightarrow\overline{B}.$
Then, we define a (not necessarily an invertible) sheaf $L_{4}$ on $B_{4}$ by
$L_{4}:=t^{w_{\gamma}({\bf f})}\bigl{(}\operatorname{det}^{-1}(I_{\overline{B}}/I_{\overline{B}}^{2})\bigr{)}\big{|}_{B_{4}}\otimes_{B_{4}}\left(\Omega_{A_{3}/k[t]}^{N}\right)\big{|}_{B_{4}},$
where $I_{\overline{B}}:=\left(f_{1},\ldots,f_{c}\right)\subset R$, and $A_{3}:=\operatorname{Spec}R[t^{1/d}]$. This $L_{4}$ will play the similar role as the $L_{\overline{B}^{(\gamma)}}$ in [NS22], and it allows us to modify the argument. A new point of this paper is that we extend the $k[t]$-schemes appearing in [NS22] to $k[t^{1/d}]$-schemes in this way, and we compare the sheaves of differentials on them.
@@ 2.1. Log pairs
A log pair $(X,\mathfrak{a})$ is a normal $\mathbb{Q}$-Gorenstein $k$-variety $X$ and an $\mathbb{R}$-ideal sheaf $\mathfrak{a}$ on $X$. Here, an $\mathbb{R}$-ideal sheaf $\mathfrak{a}$ on $X$ is a formal product $\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{r_{i}}$, where $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{s}$ are non-zero coherent ideal sheaves on $X$ and $r_{1},\ldots,r_{s}$ are positive real numbers. For a morphism $Y\to X$ and an $\mathbb{R}$-ideal sheaf $\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{r_{i}}$ on $X$, we denote by $\mathfrak{a}\mathcal{O}_{Y}$ the $\mathbb{R}$-ideal sheaf $\prod_{i=1}^{s}(\mathfrak{a}_{i}\mathcal{O}_{Y})^{r_{i}}$ on $Y$.
Let $\bigl{(}X,\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{r_{i}}\bigr{)}$ be a log pair. Let $f:X^{\prime}\to X$ be a proper birational morphism from a normal variety $X^{\prime}$ and let $E$ be a prime divisor on $X^{\prime}$. We denote by $K_{X^{\prime}/X}:=K_{X^{\prime}}-f^{*}K_{X}$ the relative canonical divisor. Then the log discrepancy of $(X,\mathfrak{a})$ at $E$ is defined as
@@
where we define $\operatorname{ord}_{E}\mathfrak{a}:=\sum_{i=1}^{s}r_{i}\operatorname{ord}_{E}\mathfrak{a}_{i}$. The image $f(E)$ is called the center of $E$ on $X$ and we denote it by $c_{X}(E)$. For a closed point $x\in X$, we define the minimal log discrepancy at $x$ as
@@ 2.2. Log triples
Let $\bigl{(}X,D,\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{r_{i}}\bigr{)}$ be a log triple. Let $f:X^{\prime}\to X$ be a proper birational morphism from a normal variety $X^{\prime}$ and let $E$ be a prime divisor on $X^{\prime}$. Then the log discrepancy of $(X,D,\mathfrak{a})$ at $E$ is defined as
@@ 2.4. Arc spaces of $k$-schemes
For $m\in\mathbb{Z}_{\geq 0}\cup\{\infty\}$, we denote by $\pi_{m}:X_{m}\to X$ the canonical truncation morphism. For $m\in\mathbb{Z}_{\geq 0}\cup\{\infty\}$ and a morphism $f:Y\to X$ of schemes of finite type over $k$, we denote by $f_{m}:Y_{m}\to X_{m}$ the morphism induced by $f$.
@@
    For an arc $\gamma\in X_{\infty}$ and an ideal sheaf $\mathfrak{a}\subset\mathcal{O}_{X}$, the order of $\mathfrak{a}$ measured by $\gamma$ is defined as follows:
    $\operatorname{ord}_{\gamma}(\mathfrak{a})=\sup\{r\in\mathbb{Z}_{\geq 0}\mid\gamma^{*}(\mathfrak{a})\subset(t^{r})\},$
    where $\gamma^{*}:\mathcal{O}_{X}\to k[[t]]$ is the induced ring homomorphism by $\gamma$.
@@
    $\operatorname{Cont}^{m}(\mathfrak{a})$ =$=\{\gamma\in X_{\infty}\mid\operatorname{ord}_{\gamma}(\mathfrak{a})=m\},$
    $\operatorname{Cont}^{\geq m}(\mathfrak{a})$ =$=\{\gamma\in X_{\infty}\mid\operatorname{ord}_{\gamma}(\mathfrak{a})\geq m\}.$
@@
    By convention, $\operatorname{codim}(C)=\infty$ if $C\cap\operatorname{Cont}^{e}\left(\operatorname{Jac}_{X}\right)=\emptyset$ for any $e\geq 0$.
@@ 2.5. Arc spaces of $k[t]$-schemes
The functor $F^{X}_{m}$ is always represented by a scheme $X_{m}$ over $k$. We shall denote by the same symbols $X_{\infty}$, $\pi_{mn}$, $\psi_{m}$, $\pi_{m}$ also for this setting. Cylinders and the contact loci $\operatorname{Cont}^{m}(\mathfrak{a})$ and $\operatorname{Cont}^{\geq m}(\mathfrak{a})$ are also defined by the same way for this setting.
@@
    By convention, $\operatorname{codim}(C)=\infty$ if $C\cap\operatorname{Cont}^{e}\bigl{(}\operatorname{Fitt}^{n}\left(\Omega_{X/k[t]}\right)\bigr{)}=\emptyset$ for any $e\geq 0$.
@@
    Let $X$ and $Y$ be $k[t]$-schemes of finite type, and let $f:X\to Y$ be a morphism over $k[t]$. Let $\gamma\in X_{\infty}$ be a $k$-arc and let $\gamma^{\prime}:=f_{\infty}(\gamma)$. Let $S$ be the torsion part of $\gamma^{*}\Omega_{X/k[t]}$. Then we define the order $\operatorname{ord}_{\gamma}(\operatorname{jac}_{f})$ of the Jacobian of $f$ at $\gamma$ as the length of the $k[[t]]$-module
    $\operatorname{Coker}\bigl{(}\gamma^{\prime*}\Omega_{Y/k[t]}\to\gamma^{*}\Omega_{X/k[t]}/S\bigr{)}.$
    In particular, if $\operatorname{ord}_{\gamma}(\operatorname{jac}_{f})<\infty$, then we have
    $\operatorname{Coker}\bigl{(}\gamma^{\prime*}\Omega_{Y/k[t]}\to\gamma^{*}\Omega_{X/k[t]}/S\bigr{)}\simeq\bigoplus_{i}k[t]/(t^{e_{i}})$
    as $k[[t]]$-modules with some positive integers $e_{i}$ satisfying $\sum_{i}e_{i}=\operatorname{ord}_{\gamma}(\operatorname{jac}_{f})$.
@@
    $\operatorname{Cont}^{e}\left(\operatorname{jac}_{f}\right):=\left\{\gamma\in X_{\infty}\ \middle|\ \operatorname{ord}_{\gamma}\left(\operatorname{jac}_{f}\right)=e\right\}$
@@
Let $n$ be a non-negative integer, and let $X$, $Y$ and $Z$ be $k[t]$-schemes. Let $f:X\to Y$ and $g:Y\to Z$ be morphisms over $k[t]$. Suppose that each irreducible component $W_{i}$ of $X$, $Y$ and $Z$ has $\dim W_{i}\geq n+1$. Let $\gamma\in X_{\infty}$ be an arc, and let $\gamma^{\prime}:=f_{\infty}(\gamma)$. Suppose that
$\operatorname{ord}_{\gamma}\left(\operatorname{Fitt}^{n}\left(\Omega_{X/k[t]}\right)\right)<\infty,\quad\operatorname{ord}_{\gamma^{\prime}}\left(\operatorname{Fitt}^{n}\left(\Omega_{Y/k[t]}\right)\right)<\infty.$
@@
$\operatorname{ord}_{\gamma}\left(\operatorname{jac}_{g\circ f}\right)=\operatorname{ord}_{\gamma}\left(\operatorname{jac}_{f}\right)+\operatorname{ord}_{\gamma^{\prime}}\left(\operatorname{jac}_{g}\right).$
@@
Let $n$ and $N$ be non-negative integers with $n\leq N$. Set $c:=N-n$. Let $f_{1},\ldots,f_{c}\in k[x_{1},\ldots,x_{N}][t]$, and let $I_{X}=(f_{1},\ldots,f_{c})\subset k[x_{1},\ldots,x_{N}][t]$ be the ideal generated by them. We set
@@
  • $C\cap X_{\infty}\cap\operatorname{Cont}^{e}\left(\operatorname{Fitt}^{n}\left(\Omega_{X/k[t]}\right)\right)\not=\emptyset$

### 30 Apr 2026 v2

@@ Abstract.
We prove the precise inversion of adjunction formula for finite linear group quotients of complete intersection varieties defined by semi-invariant equations. As an application, we prove the semi-continuity of minimal log discrepancies for them. These results extend the results in our first paper, where we prove the same results for finite linear group quotients of complete intersection varieties defined by “invariant equations”.
@@ 1. Introduction
The main purpose of this paper is to relax the assumptions of “invariant equations” in the results in [NS22] and extend them to the case of “semi-invariant equations”. For a finite subgroup $G\subset{\rm GL}_{N}(k)$, we say that $f\in k[x_{1},\ldots,x_{N}]$ is $G$-semi-invariant if for any $\gamma\in G$, there exists $s\in k^{\times}$ such that $\gamma(f)=sf$. In this paper, we show that the LSC conjecture is true for a finite linear group quotient of a complete intersection variety defined by semi-invariant equations.
The advantage of being able to handle semi-invariant equations is that it enables us to treat a class containing 3-dimensional terminal singularities. According to the classification by Reid and Mori ([Rei80, Rei81, Mor85]), 3-dimensional terminal singularities are known to be finite quotients of smooth points or cDV points. As seen in Mori’s classification list ([Mor85]), semi-invariant equations indeed appear alongside invariant equations.
Theorem 1.2 (== Theorem 7.4).
@@
Let $N$ be a positive integer. Set $R:=k[x_{1},\ldots,x_{N}]$. Let $d$ be a positive integer, and let $\xi\in k$ be a primitive $d$-th root of unity. Suppose that a finite group $G\subset\operatorname{GL}_{N}(k)$ of order $d$ acts on $\overline{A}=\operatorname{Spec}R$. Let $\overline{B}\subset\overline{A}$ be a complete intersection variety defined by $G$-semi-invariant elements $f_{1},\ldots,f_{c}\in(x_{1},\ldots,x_{N})\subset R$. We denote by $B:=\overline{B}/G$ its quotient variety. Since $G$ is a finite group, each $\gamma$ can be diagonalized with respect to a new basis $x_{1}^{(\gamma)},\ldots,x_{N}^{(\gamma)}$ consisting of $k$-linear combinations of the original basis $x_{1},\ldots,x_{N}$. Let $\operatorname{diag}\left(\xi^{e_{1}},\ldots,\xi^{e_{N}}\right)$ be the diagonal matrix with $0\leq e_{i}\leq d-1$. Then we define a $k$-algebra homomorphism ${\lambda}^{*}_{\gamma}:R\to R[t^{1/d}]$ as follows:
@@
In [NS22], we study the minimal log discrepancies of $B$ via the arc space $\overline{B}^{(\gamma)}_{\infty}$ by applying Denef and Loeser’s theory [DL02]. Here, $\overline{B}^{(\gamma)}_{\infty}$ is the arc space of $\overline{B}^{(\gamma)}$ as a $k[t]$-scheme (Subsection 2.4), which is essentially the corresponding Greenberg scheme (see Remark 2.7 for details).
However, if $f_{i}$ is just assumed to be $G$-semi-invariant, $\lambda^{*}_{\gamma}(f_{i})$ is not necessarily an element of $R[t]$. Thus, the first difficulty is how to redefine $\overline{B}^{(\gamma)}$. For each $i$ and $\gamma\in G$, we define $w_{\gamma}(f_{i}):=\frac{a}{d}$ when $\gamma(f_{i})=\xi^{a}f_{i}$ with $0\leq a\leq d-1$. Then, we have $\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\in R[t]$. By using this invariant $w_{\gamma}$, we change the definition of $\overline{B}^{(\gamma)}$ as follows:
@@
Theorem 1.6 (== Theorem 5.3).
Let n:=Ncn:=N-c. Let $x\in B$ be the image of the origin of $\overline{A}=\mathbb{A}_{k}^{N}$. Let $\mathfrak{a}\subset\mathcal{O}_{B}$ be a non-zero ideal sheaf and $\tau$ a positive real number. Then
$\operatorname{mld}_{x}(B,\mathfrak{a}^{\tau})$ $=\inf_{\gamma\in G,\ b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C_{\gamma,b_{1},b_{2}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\tau b_{1}\right\}$
$=\inf_{\gamma\in G,\ b_{1},b_{2}\in\mathbb{Z}_{\geq 0}}\left\{\operatorname{codim}\left(C^{\prime}_{\gamma,b_{1},b_{2}}\right)+\operatorname{age}(\gamma)-w_{\gamma}({\bf f})-b_{2}-\tau b_{1}\right\}$
@@
By the theorem of Ein, Mustaţă, and Yasuda ([EMY03]), the minimal log discrepancy $\operatorname{mld}_{x}(B,\mathfrak{a}^{\tau})$ is described by the codimensions of the corresponding contact loci in the arc space $B_{\infty}$ of $B$. Therefore, in order to prove Theorem 1.6, we need to compare the codimension of a cylinder $C\subset B_{\infty}$ and the codimension of the corresponding cylinder $C^{\prime}\subset\overline{B}^{(\gamma)}_{\infty}$. By the same argument in [NS22] (using the theory of Denef and Loeser in [DL02]), the difference can be measured by comparing $\Omega_{\overline{B}^{(\gamma)}/k[t]}^{n}$ and $\omega_{B}$ via the morphism $\overline{B}^{(\gamma)}\to B$.
@@
$L_{\overline{B}^{(\gamma)}}:=\bigl(\operatorname{det}^{-1}(I_{B}/I_{B}^{2})\bigr)\big|_{\overline{B}^{(\gamma)}}\otimes_{\mathcal{O}_{\overline{B}^{(\gamma)}}}\Bigl(\Omega_{\overline{A}^{(\gamma)}/k[t]}^{N}\Bigr)\big|_{\overline{B}^{(\gamma)}},$
@@
$\overline{B}^{(\gamma)}_{1/d}:=\overline{B}^{(\gamma)}\times_{\operatorname{Spec}k[t]}\operatorname{Spec}k[t^{1/d}]\longrightarrow\overline{B}.$
Then, we define a (not necessarily an invertible) sheaf $L_{\overline{B}^{(\gamma)}_{1/d}}$ on $\overline{B}^{(\gamma)}_{1/d}$ by
$L_{\overline{B}^{(\gamma)}_{1/d}}:=t^{w_{\gamma}({\bf f})}\bigl(\operatorname{det}^{-1}(I_{\overline{B}}/I_{\overline{B}}^{2})\bigr)\big|_{\overline{B}^{(\gamma)}_{1/d}}\otimes_{\overline{B}^{(\gamma)}_{1/d}}\left(\Omega_{\overline{A}^{(\gamma)}_{1/d}/k[t]}^{N}\right)\big|_{\overline{B}^{(\gamma)}_{1/d}}$,
where $I_{\overline{B}}:=\left(f_{1},\ldots,f_{c}\right)\subset R$, and $\overline{A}^{(\gamma)}_{1/d}:=\operatorname{Spec}R[t^{1/d}]$. This $L_{\overline{B}^{(\gamma)}_{1/d}}$ will play the similar role as the $L_{\overline{B}^{(\gamma)}}$ in [NS22], and it allows us to modify the argument. A new point of this paper is that we extend the $k[t]$-schemes appearing in [NS22] to $k[t^{1/d}]$-schemes in this way, and we compare the sheaves of differentials on them.
@@
The klt assumption on $Y$ and $D$ in Theorem 1.4 is used to prove $b_{3}<\infty$ in the proof of Theorem 6.4. We cite [NS22]*Claim 5.2 for this step, which is essentially based on the following two results: the result by Hacon and McKernan [HM07], which establishes the rational chain connectedness of fibers of resolutions of klt singularities; and the result by Graber, Harris, and Starr [GHS03], which asserts the existence of a section for a morphism with rational chain connected fibers. We emphasize that this assumption cannot be dropped, as [NS4] constructs a counterexample in the non-klt setting.
@@ 2.1. Log pairs
A log pair $(X,\mathfrak{a})$ is a normal $\mathbb{Q}$-Gorenstein $k$-variety $X$ and an $\mathbb{R}$-ideal sheaf $\mathfrak{a}$ on $X$. Here, an $\mathbb{R}$-ideal sheaf $\mathfrak{a}$ on $X$ is a formal product $\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{\tau_{i}}$, where $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{s}$ are non-zero coherent ideal sheaves on $X$ and $\tau_{1},\ldots,\tau_{s}$ are positive real numbers. For a morphism $Y\to X$ and an $\mathbb{R}$-ideal sheaf $\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{\tau_{i}}$ on $X$, we denote by $\mathfrak{a}\mathcal{O}_{Y}$ the $\mathbb{R}$-ideal sheaf $\prod_{i=1}^{s}(\mathfrak{a}_{i}\mathcal{O}_{Y})^{\tau_{i}}$ on $Y$.
Let $\bigl(X,\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{\tau_{i}}\bigr)$ be a log pair. Let $f:X^{\prime}\to X$ be a proper birational morphism from a normal variety $X^{\prime}$ and let $E$ be a prime divisor on $X^{\prime}$. We denote by $K_{X^{\prime}/X}:=K_{X^{\prime}}-f^{*}K_{X}$ the relative canonical divisor. Then the log discrepancy of $(X,\mathfrak{a})$ at $E$ is defined as
@@
where we define $\operatorname{ord}_{E}\mathfrak{a}:=\sum_{i=1}^{s}\tau_{i}\operatorname{ord}_{E}\mathfrak{a}_{i}$. The image $f(E)$ is called the center of $E$ on $X$ and we denote it by $c_{X}(E)$. For a closed point $x\in X$, we define the minimal log discrepancy at $x$ as
@@ 2.2. Log triples
Let $\bigl(X,D,\mathfrak{a}=\prod_{i=1}^{s}\mathfrak{a}_{i}^{\tau_{i}}\bigr)$ be a log triple. Let $f:X^{\prime}\to X$ be a proper birational morphism from a normal variety $X^{\prime}$ and let $E$ be a prime divisor on $X^{\prime}$. Then the log discrepancy of $(X,D,\mathfrak{a})$ at $E$ is defined as
@@ 2.4. Arc spaces of $k$-schemes
For $m\in\mathbb{Z}_{\geq 0}\cup\{\infty\}$ and a morphism $f:Y\to X$ of schemes of finite type over $k$, we denote by $f_{m}:Y_{m}\to X_{m}$ the morphism induced by $f$.
@@
    For an arc $\alpha\in X_{\infty}$ represented by a morphism $\alpha:\operatorname{Spec}K[[t]]\to X$ (where $K$ is a field extension of $k$) and an ideal sheaf $\mathfrak{a}\subset\mathcal{O}_{X}$, the order of $\mathfrak{a}$ measured by $\alpha$ is defined as follows:
    $\operatorname{ord}_{\alpha}(\mathfrak{a})=\sup\{r\in\mathbb{Z}_{\geq 0}\mid\alpha^{*}(\mathfrak{a})\subset(t^{r})\}\in\mathbb{Z}_{\geq 0}\cup\{\infty\},$
    where $\alpha^{*}(\mathfrak{a})$ denotes the ideal of $K[[t]]$ generated by the image of $\mathfrak{a}$.
@@
    $\operatorname{Cont}^{m}(\mathfrak{a})$ $=\{\alpha\in X_{\infty}\mid\operatorname{ord}_{\alpha}(\mathfrak{a})=m\},$
    $\operatorname{Cont}^{\geq m}(\mathfrak{a})$ $=\{\alpha\in X_{\infty}\mid\operatorname{ord}_{\alpha}(\mathfrak{a})\geq m\}.$
@@
    By convention, $\operatorname{codim}(C)=\infty$ if $C\cap\operatorname{Cont}^{e}\left(\operatorname{Jac}_{X}\right)=\emptyset$ for every $e\geq 0$.
Remark 2.5.
In what follows, we only consider $k$-arcs (i.e., we assume $K=k$). The theory of arc spaces employed in this paper primarily focuses on cylinders and their codimensions. Since the codimension of a cylinder in $X_{\infty}$ is defined at the level of finite jet schemes and $k$ is an algebraically closed field, restricting our attention to $k$-arcs is sufficient for our purposes.
@@ 2.5. Arc spaces of $k[t]$-schemes
The functor $F^{X}_{m}$ is always represented by a scheme $X_{m}$ over $k$. We use the same notation $X_{\infty}$, $\psi_{m}$, and $\pi_{mn}$ in this setting as well:
[…]
We also define a subset $C\subset X_{\infty}$ to be a cylinder if it is of the form $C=\psi_{m}^{-1}(S)$ for some $m\geq 0$ and a constructible subset $S\subset X_{m}$. The contact loci $\operatorname{Cont}^{m}(\mathfrak{a})$ and $\operatorname{Cont}^{\geq m}(\mathfrak{a})$ are defined in the same way as in Definition 2.3.
@@
Remark 2.7.
The schemes $X_{m}$ and $X_{\infty}$ coincide with the Greenberg schemes $\operatorname{Gr}_{m}(\mathfrak{X})$ and $\operatorname{Gr}_{\infty}(\mathfrak{X})$ (defined in [Seb04] and [CLNS]) for the formal scheme $\mathfrak{X}$ over $k[[t]]$ associated to the $k[t]$-scheme $X$. See [NS22]*Remark 2.14 for details.
@@
    By convention, $\operatorname{codim}(C)=\infty$ if $C\cap\operatorname{Cont}^{e}\bigl(\operatorname{Fitt}^{n}\left(\Omega_{X/k[t]}\right)\bigr)=\emptyset$ for every $e\geq 0$.
@@
    Let $X$ and $Y$ be $k[t]$-schemes of finite type, and let $f:X\to Y$ be a morphism over $k[t]$. Let $\alpha\in X_{\infty}$ be a $k$-arc and let $\alpha^{\prime}:=f_{\infty}(\alpha)$. Let $S$ be the torsion part of $\alpha^{*}\Omega_{X/k[t]}$. Then we define the order $\operatorname{ord}_{\alpha}(\operatorname{jac}_{f})$ of the Jacobian of $f$ at $\alpha$ as the length of the $k[[t]]$-module
    $\operatorname{Coker}\bigl(\alpha^{\prime*}\Omega_{Y/k[t]}\to\alpha^{*}\Omega_{X/k[t]}/S\bigr).$
    In particular, if $\operatorname{ord}_{\alpha}(\operatorname{jac}_{f})<\infty$, then we have
    $\operatorname{Coker}\bigl(\alpha^{\prime*}\Omega_{Y/k[t]}\to\alpha^{*}\Omega_{X/k[t]}/S\bigr)\simeq\bigoplus_{i}k[t]/(t^{e_{i}})$
    as $k[[t]]$-modules with some positive integers $e_{i}$ satisfying $\sum_{i}e_{i}=\operatorname{ord}_{\alpha}(\operatorname{jac}_{f})$.
@@
    $\operatorname{Cont}^{e}\left(\operatorname{jac}_{f}\right):=\left\{\alpha\in X_{\infty}\ \middle|\ \operatorname{ord}_{\alpha}\left(\operatorname{jac}_{f}\right)=e\right\}$
@@
Let $n$ be a non-negative integer, and let $X$, $Y$ and $Z$ be $k[t]$-schemes. Let $f:X\to Y$ and $g:Y\to Z$ be morphisms over $k[t]$. Suppose that each irreducible component $W_{i}$ of $X$, $Y$ and $Z$ has $\dim W_{i}\geq n+1$. Let $\alpha\in X_{\infty}$ be an arc, and let $\alpha^{\prime}:=f_{\infty}(\alpha)$. Suppose that
$\operatorname{ord}_{\alpha}\left(\operatorname{Fitt}^{n}\left(\Omega_{X/k[t]}\right)\right)<\infty,\quad\operatorname{ord}_{\alpha^{\prime}}\left(\operatorname{Fitt}^{n}\left(\Omega_{Y/k[t]}\right)\right)<\infty.$
@@
$\operatorname{ord}_{\alpha}\left(\operatorname{jac}_{g\circ f}\right)=\operatorname{ord}_{\alpha}\left(\operatorname{jac}_{f}\right)+\operatorname{ord}_{\alpha^{\prime}}\left(\operatorname{jac}_{g}\right).$
@@
Let $n$ and $N$ be non-negative integers with $n\leq N$. Set c:=Nnc:=N-n. Let $f_{1},\ldots,f_{c}\in k[x_{1},\ldots,x_{N}][t]$, and let $I_{X}=(f_{1},\ldots,f_{c})\subset k[x_{1},\ldots,x_{N}][t]$ be the ideal generated by them. We set
@@
  • $C\cap X_{\infty}\cap\operatorname{Cont}^{e}\left(\operatorname{Fitt}^{n}\left(\Omega_{X/k[t]}\right)\right)\not$$\emptyset$